OpenStudy (anonymous):

Identify the correct x intercept and y intercept of the equation x - 2y = 8. Answer x intercept = (-8, 0); y intercept = (0, -4) x intercept = (8, 0); y intercept = (0, -4) x intercept = (8, 0); y intercept = (0, 4) x intercept = (-2, 0); y intercept = (0, 8)

OpenStudy (callisto):

For x-int., put y=0 into the equation and solve x x - 2y = 8 x-2(0) = 8 x=8 So, x-int is at (8, 0) For y-int., put x=0 into the equation and solve y Can you find the y-int?

OpenStudy (anonymous):

nope

OpenStudy (callisto):

Put x=0 into x - 2y = 8 0 - 2y =8 -2y=8 Divide both sides by coefficient of y, what do you get?

OpenStudy (anonymous):

idk

OpenStudy (callisto):

What is the coefficient of y?

OpenStudy (anonymous):

idk how bout yu just give me the answer